OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10

OpenStudy (anonymous):

http://www.algebrahelp.com/calculators/equation/ type it here. and u will get the answer

OpenStudy (anonymous):

for positive charge we can write -5(x+1)=10 =>-5x-5=10 =>-5x=10+5 =>-5x=15 =>x=-3 for negative charge we can write -.-5(x+1)=10 =>5(x+1)=10 =>5x+5=10 =>5x=10-5 =>x=1